OpenStudy (anonymous):

How do we rewrite -2y+5= -6x so that y is a function of x in function notation? I will give metal to who ever answers first!

OpenStudy (anonymous):

...

OpenStudy (triciaal):

y=3x +5/2

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

cause i got the same thing except i actually divided 5 by 2 y=3x+2.5

OpenStudy (anonymous):

As a general rule, do your additions and subtractions first and then your divisions or multiplications. a y + b = c x + d ay = c x + d - b y = [ c x + d - b]/a